OpenStudy (anonymous):

What happens when you multiply a negative number by another negative number? Is it still negative or does it become positive?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

same with division

OpenStudy (amistre64):

the rule of thumb; whenever you have an even number of same signs; go positive

OpenStudy (anonymous):

So... -6(x+-6) is -6x +-6?

OpenStudy (amistre64):

more like -6(x+-6) = -6x +--6(6) ^ -- means +; just stab one with the other to make a + sign

OpenStudy (anonymous):

So it actually equals 36.

OpenStudy (amistre64):

-6(x+-6) -6x +--6(6) -6x ++36 -6x + 36

OpenStudy (amistre64):

so yes :)

OpenStudy (anonymous):

Wow....ok thanks x)

OpenStudy (anonymous):

good job jess! (and amistre)

OpenStudy (amistre64):

.... i have to confess, ive actually done these things before lol

OpenStudy (amistre64):

i still do it on my calculus tests just to make sure i dont confuse me signs

OpenStudy (anonymous):

I occasionally still count on my fingers :)

OpenStudy (anonymous):

we can show -*- = + by considering numbers and brackets 5 - 3(4 - 1) working brackets first this = 5 - 3*3 = -4 using the distributive law -4 can only be obtained if -*- = + 5 - 3*4 -3*-1 = 5-12+3 = -4

OpenStudy (amistre64):

adding with my fingers is easy; its subtraction that gets painful ;)

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

What is *3? I don't understand what you did.

OpenStudy (anonymous):

*3 is "times 3"